solve the following quadratic equations using the completing the squares.
A.) x2 + 6x - 12 = 0
B.) x2 + 2x = 24

Answer by lwsshak3(11628)

solve the following quadratic equations using the completing the squares.
A.) x2 + 6x - 12 = 0
(x^2+6x+9)-9-12=0
(x+3)^2-21=0
(x+3)^2=21
take sqrt of both sides
x+3=±√21
x=-3+√21
or
x=-3-√21
..
B.) x2 + 2x = 24
(x^2+2x+1)=24+1
(x+1)^2=25
take sqrt of both sides
x+1=±√25
x=-1±5
x=-6
or
x=4